The word problem is: Steve is six years older than Sarah is now. In ten years he will be twice her age. Find their ages.
I figured s= steve's age now, and x=sarah's age now. so s+10=Steve in 10 years. and 2x+10=Sarah in 10 years. This is where i start to get confused.
s+10=2(x+10). so does that make 6+x+10=2(x+10)? If thats true then when i finisht the math why do i get x=-4 or -x=4. can you help please?

Answer by stanbon(75887)

The word problem is: Steve is six years older than Sarah is now. In ten years he will be twice her age. Find their ages.
----
Equations:
t = s+6
(t+10) = 2(s+10)
-----
Substitute for "t" and solve for "s":
(s+6+10) = 2s+20
s + 16 = 2s+20
s = -4 yrs old (Sarah will not be born till 4 years pass)
------
t = s+6 = -4+6 = 2 (Steve is 2 years old now
======================================================
Some problems just don't make sense.
This is one of them.
=========================================
Cheers,
Stan H.